1.What is the price (per $1,000 of face value) of a seven-year 8 percent coupon bond? Assume that the discount rate is 6 percen

2.Imagine the bond described in the previous question has a market price of $1,300. Using Excel, find the yield to maturity of

(Preferably for #2 using Goal Seek tool) Please show steps if able
Answer #1

Answer:

1) Given

Face value F=$1000

Coupon rate =8%

Coupon C=8%*1000=$80

N= 7 years

Discount rate R=6%

Bond Price P= C*(1-(1+R)^-N)/R + F/(1+R)^N

P=80*(1-(1+6%)^-7)/6% + 1000/(1+6%)^7=$1111.65

2)

Given Bond price P=$1300

Face value F=$1000

Coupon rate =8%

Coupon C=8%*1000=$80

N= 7 years

Let Yield to maturity is R

So

Bond Price P= C*(1-(1+R)^-N)/R + F/(1+R)^N

1300=80*(1-(1+R)^-7)/R + 1000/(1+R)^7

Solving for R we get

R=3.16%